OpenStudy (anonymous):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank. 1 over 3π(x - 7)^2 (x - 3) What does the factor π (x - 7)^2 represent? A. The total surface area of the tank B. The area of the curved sides of the tank C. The area of the circular base of the tank D. The area of the base of about three such tanks

OpenStudy (anonymous):

Help PLEASE!!!!!!!!!!!!!

OpenStudy (anonymous):

The volume of a cone is \(\dfrac{1}{3}\pi r^2h\). The factor \(\pi r^2\) is the area of the circular base.

OpenStudy (anonymous):

Thank you so much!!! How do I give you a medal?